OpenStudy (brucebaner):

Complete the standard form of the equation . x + y = 9

OpenStudy (brucebaner):

Complete the standard form of the equation . \[-4/3 y=3-x\]

OpenStudy (brucebaner):

@abb0t

OpenStudy (brucebaner):

@girlover

OpenStudy (anonymous):

-4/3y or (-4/3)y ?

OpenStudy (brucebaner):

-4/3y

OpenStudy (anonymous):

-4/3y = 3-x -4 = (3-x)3y 3y = -4/(3-x) y = -4 / [3(3-x)]

OpenStudy (brucebaner):

plz wait